What is the charge of a body whose proton is 4 x 10 15 and 5 x 10 15 ?
1
Expert's answer
2020-12-14T07:16:27-0500

The charge of one proton is


"e = 1.6\times 10^{-19}C"

If a body contains "N_1 = 4\times 10^{15}" or "N_2 = 5\times 10^{15}" uncompensated protons, then its charge is (respectively):


"Q_1 = eN_1 = 1.6\times 10^{-19}C\cdot 4\times 10^{15} = 6.4\times 10^{-4}C\\
Q_2 = eN_2 = 1.6\times 10^{-19}C\cdot 5\times 10^{15} = 8\times 10^{-4}C"

Answer.

"Q_1 = 6.4\times 10^{-4}C\\
Q_2 = 8\times 10^{-4}C"